A cylinder with high rigidity, diameter inside it is 25 mm containing a column of water 650 mm
long. Calculate the column length if a force of 2.5 kN is applied to its end by a frictionless plunger?
Assume there is no leakage.
1
Expert's answer
2020-06-11T15:10:36-0400

K =−∆p ∆V/V K =2.5×10^3N.

∆ p=P −P   (P =0.0 because no applied forces in first case)

P =Force/ Area=2.5×10^3/ π 4×0.025

= 5 ×10Pa.

 →∆p=3.25×10−0 = 3.25×10Pa. 

∆V V=V−V V =L×A−0.65×A

0.65×A

=L−0.65 0.5 → K=−∆ ∆ →2×10^3

=(−3.25*10^6 × 0.65)/(L-0.65)

→L=0.637m.